a machine of mass 100 kg is supported on openings of total stiffness 800 kn/m and has a rotating unbalanced element which results in a disturbing force of 400 n at a speed of 3000 r.p.m. assum-ing the damping ratio as 0.25, determine : 1. the amplitude of vibrations due to unbalance ; and 2. the transmitted force.
1
Expert's answer
2021-09-21T02:06:40-0400

"m = 100\textsf{ kg}\\
s = 800\textsf{ kN/m}L\\
F = 400\textsf{ N}\\
\omega = 3000\textsf{ rpm}= 314.2\textsf{ rad/s}\\
\textsf{(using 1 rev = 2π rad)}\\
x_2 = x_1/4"


We know that angular speed or natural circular frequency of free vibrations;

"\omega_n= \sqrt\dfrac sm= \sqrt\dfrac{800×10³}{100}= 89.44 \textsf{ rad/s}"


For the amplitude;

"\log_e \dfrac{x_1}{x_2} = a× \dfrac{2π}{\sqrt{(\omega_n)²-a²}}"


"\log_e4 = \dfrac{2πa}{\sqrt{(89.44)²-a²}}"


"a =\dfrac{ 2236 log(2)}{(25 \sqrt{(π^2 + log4)}}"


"a=19.27"


"c= a × 2m = 19.27 × 2 × 10= 385.4\textsf{ N/m/s}"


"x_o = \dfrac Fs = \dfrac{400}{800× 10³}= 0.0005m"


Amplitude of the forced vibration;


"x_{\textsf{max}} = \dfrac{x_o}{\sqrt{\dfrac{c²\omega²}{s²}+[1-\dfrac{\omega²}{\omega_n²}]²}}"


"= \dfrac{0.0005}{\sqrt{\dfrac{385.4²×314.2²}{(800×10³)²}+[1-\dfrac{314.2²}{89.44²}]²}}"


"= 0.00004408m = 0.0044mm"


[0.044mm; 35.2 N]